OpenStudy (anonymous):

Evaluate: [(21 + 6) − 32] ÷ 9 • 2 4 4.6 1 52

OpenStudy (diyadiya):

Use PEMDAS

Parth (parthkohli):

Yes^^

Parth (parthkohli):

@Diyadiya No, first solve the brackets, then use PEMDAS/BODMAS(whatever you call it)

OpenStudy (anonymous):

use a calculator

OpenStudy (diyadiya):

Parentheses, Exponents, Multiplication and Division, and Addition and Subtraction

OpenStudy (diyadiya):

@staceyyyy1 did you understand ?

OpenStudy (anonymous):

No. Not At All.

OpenStudy (anonymous):

Okay..

OpenStudy (diyadiya):

what do you get now ?

OpenStudy (anonymous):

27

Parth (parthkohli):

@Diyadiya We have to solve the parentheses first if they are given, then use PEMDAS.

OpenStudy (anonymous):

I think you won't get any of the options given.

OpenStudy (diyadiya):

Oh yes parth is right !! I meant to solve the Parentheses 21+6=27 now 27-32=-5 9. 2=18 -5÷9.2 -5/18

OpenStudy (anonymous):

4.6